Price Quantity demanded

R1 100

R2 80

R3 60

R4 40

R5 20

What is the price elasticity of demand when the price changes from R4 per ice cream cone to R5 per ice cream cone? (Use the formula for arc elasticity.)


1
Expert's answer
2021-07-28T14:23:03-0400

When the price changes from R4 per ice cream cone to R5 per ice cream cone, then the price elasticity of demand is:

"Ed = \frac{20 - 40} {5 - 4} ×\frac{5 + 4} {20 + 40} = -3."

So, the demand is elastic at this interval.